OpenStudy (anonymous):

Solve: x2 – 16x + 15 = 0

OpenStudy (saifoo.khan):

15 or 1

OpenStudy (anonymous):

try \[(x-15)(x-1)=\] \[x-15=0\] \[x=15\] or \[x-1=0\] \[x=1\]